OpenStudy (anonymous):

Three liquids of densities d, 2d, and 3d are mixed in equal volumes. Then the density of the mixture is: (a) d (b) 2d (c) 3d (d) 5d @Mathematics

OpenStudy (anonymous):

2d

OpenStudy (anonymous):

Can you explain how?

OpenStudy (anonymous):

d=m/V m1=dV m2=2dV m3=3dV then d(mixture)=(m1+m2+m3)/(V+V+V)=(6dV)/(3V)=2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

the answer must be 2d not 2:)